# Open Vocabulary Monocular 3D Object Detection

### v1

@@ 5.1 Experimental Setup
The experiments are conducted using the Omni3D [5] dataset, the largest publicly available dataset for monocular 3D object detection across both indoor and outdoor scenes. Omni3D is repurposed and combined from six established sources including SUN RGB-D [51], ARKitScenes [2], Objectron [1], Hypersim [45], nuScenes [6], and KITTI [15]. It comprises a substantial 234k images, 3 million labeled 3D bounding boxes, and covers 98 distinct object categories.
During training, we use the 50 categories Cube R-CNN [5] trained on as the base categories for model training. For evaluation, we select 22 categories from the remaining classes as novel categories. These categories are chosen based on two criteria: the number of test instances and the precision of category naming. To facilitate a detailed assessment of zero-shot generalization capabilities, we further divide these categories into easy and hard subsets according to object visibility. The easy categories include board, printer, painting, microwave, tray, podium, cart, and tram, while the remaining categories constitute the hard subset.
| Methods | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ |
| G-DINO-3D (FT) | 6.77 | 4.43 | 6.21 | 4.92 | 1.12 | 4.77 | 4.24 |
| G-DINO-3D (FZ) | 19.22 | 3.54 | 5.74 | 2.71 | 0.07 | 4.59 | 2.96 |
| OVMono3D-GEO | 18.72 | 9.03 | 14.88 | 7.60 | 0.43 | 12.55 | 7.01 |
| OVMono3D-LIFT | 21.59 | 16.05 | 24.27 | 14.97 | 2.40 | 20.58 | 13.47 |
Evaluations.
We report a mean AP across all 50 base categories and 22 novel categories in Omni3D and over a range of IoU3D thresholds . Unless explicitly stated otherwise, all the evaluations are under our target-aware evaluation protocols (Sec. 3.2).
Besides the standard IoU metrics, we also report the Normalized Hungarian Distance (NHD) [3] which provides a more precise evaluation. NHD computes the normalized distance between predicted and ground-truth cube corners using the Hungarian matching algorithm. NHD is computed for predictions that have a high IoU with the ground-truth 2D bounding boxes. Therefore, NHD values for different methods with varying 2D inputs are not directly comparable. To disentangle the error contributions of different predicted attributes to the overall 3D bounding box prediction, we report disentangled NHD for the position (XY), depth, dimensions, and pose. For each attribute, we compute the NHD of a disentangled 3D bounding box that uses the predicted value for that attribute while setting the other attributes to their ground-truth values, comparing it with the ground-truth cube.
@@
Our implementation is based on Pytorch3D [43] and Detectron2 [56]. We utilize DINOv2 [38] as the image feature encoder, freezing all parameters during training. All experiments are conducted on four NVIDIA A100 GPUs. The models are trained for 116k steps with a batch size of 64. We employ SGD with an initial learning rate of 0.012, which decays by a factor of 10 at 60% and 80% of the training duration. During training, image augmentations including random horizontal flipping and resizing are applied.
| AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | |
| Cube R-CNN [5] | 32.31 | 26.78 | 35.71 | 28.72 | 10.48 |
| OVMono3D-LIFT | |||||
| + Grounding DINO [32] | 21.96 | 24.77 | 33.23 | 26.33 | 9.38 |
| + Cube R-CNN [5] | 32.45 | 26.44 | 35.30 | 28.40 | 10.26 |
| + Ground Truth | 100 | 29.07 | 39.74 | 30.38 | 10.39 |
@@ 5.2 Model Performance
Fig. 5 shows qualitative results on the Omni3D test set. In comparison with Cube R-CNN [5], OVMono3D-LIFT detects not only objects of base categories, but also novel categories that are unseen at the training time.
Novel Category Performance.
Tab. 1 shows that OVMono3D-LIFT outperforms the baseline methods on novel categories by a large margin. OVMono3D-GEO demonstrates lower performance due to its sensitivity to occlusions and inaccuracies in metric depth map estimation. G-DINO-3D (FZ) freezes the 2D detection parameters, resulting in object features that are not well-suited for 3D detection. Conversely, G-DINO-3D (FT) fine-tunes the model, degrading its original OV detection capabilities, since monocular 3D detection lacks the extensive detection and grounding annotations available in 2D. Therefore, these approaches are suboptimal. In contrast, OVMono3D-LIFT effectively leverages 2D data-driven priors, including OV 2D detectors and self-supervised 2D features, by decoupling the OVMono3D task into OV 2D detection and 2D-to-3D unprojection. This approach mitigates issues related to low-quality and scarce annotations in 3D, leading to superior performance.
@@
Tab. 2 compares OVMono3D-LIFT with Cube R-CNN on base categories using different 2D bounding box inputs. When using the OV 2D detector’s predictions as input, OVMono3D-LIFT shows lower performance due to the OV 2D detector’s lower zero-shot performance on the Omni3D test set. In contrast, when using Cube R-CNN’s 2D predictions, OVMono3D-LIFT achieves comparable performance to Cube R-CNN. Notably, Cube R-CNN utilizes per-category priors for object dimensions and depth, highlighting the superior performance of OVMono3D-LIFT without relying on such priors. Additionally, we provide results with ground-truth 2D inputs as a reference.
| Original Evaluation | Target-Aware Evaluation | |||||||
| Models | AP | AP | AR | AR | ||||
| Cube R-CNN (Base) | 27.65 | 23.68 | 42.35 | 37.56 | 32.31+4.66 | 26.78+3.10 | 42.35+0.00 | 37.56+0.00 |
| OVMono3D-LIFT (Base) | 12.27 | 17.22 | 27.19 | 33.45 | 21.96+9.69 | 24.77+7.55 | 40.13+12.94 | 47.28+13.83 |
| OVMono3D-LIFT (Novel) | 7.01 | 4.22 | 17.86 | 15.58 | 21.59+14.58 | 16.05+11.83 | 37.88+20.02 | 36.85+21.27 |
| Methods | NHD | XY | Depth | Size | Pose |
| OVMono3D-GEO | 5.47 | 2.23 | 4.30 | 2.34 | 2.04 |
| OVMono3D-LIFT | 5.72 | 2.28 | 4.73 | 1.05 | 0.55 |
@@
| Ground Truth | 100 | 17.21 | 25.71 | 16.30 | 2.79 |
| YOLO-World | 20.14 | 12.82 | 20.25 | 11.91 | 1.63 |
| Grounding DINO | 21.59 | 16.05 | 24.27 | 14.97 | 2.40 |
5.3 Analysis
Evaluation Protocol.
Tab. 3 shows that our target-aware evaluation leads to increased metrics on both base and novel category settings. Even for the original Cube R-CNN, our method demonstrates improvements in both AP and AP, indicating its effectiveness in eliminating the adverse effects of missing annotations. On novel categories, our target-aware evaluation shows considerable metric value gains. This improvement is particularly important because novel categories are often infrequent and smaller in size, making them harder for the model to detect when prompted with an extensive list of possible categories.
Disentangled Metrics.
Tab. 4 exhibits the overall and disentangled NHD for different attribute predictions. For these two methods, object depth prediction consistently contributes the most error to the overall prediction. This indicates that object depth estimation is the primary bottleneck of our task. Furthermore, the size and pose predictions in OVMono3D-GEO exhibit larger errors, indicating that geometry-based methods are not effective for these predictions. This underscores the necessity of developing learning-based methods.
2D Bounding Box Input.
Tab. 5 evaluates the impact of 2D detectors on both 2D and 3D detection performance for novel categories. We test two state-of-the-art 2D detectors: YOLO-World [12] and Grounding DINO [32], alongside ground-truth 2D bounding boxes as reference. Grounding DINO consistently outperforms YOLO-World across all evaluation metrics. Notably, the 3D lifting results from Grounding DINO’s zero-shot 2D detection are only 1% lower than the performance with ground-truth 2D input in terms of AP, demonstrating that Grounding DINO is a preferable off-the-shelf 2D detector for OVMono3D.
Pre-trained Backbone Selection.
Tab. 6 shows the impact of pretrained feature extractors on 3D detection performance for novel categories. DINOv2 achieves the best performance across all evaluation metrics. This result underscores the effectiveness of DINOv2’s representations for the 3D detection task. Our findings align with recent studies that highlight DINOv2’s strong zero-shot capabilities in understanding depth, multi-view correspondences, and relative pose [27, 17, 62], demonstrating that DINOv2’s features are highly suitable for this task.
| Feature Extractor | Arch. | Supervision | Dataset | AP↑ | AP↑ | AP↑ | AP↑ |
| MAE [21] | ViT-B/16 | SSL | ImageNet1k | 7.72 | 11.95 | 6.43 | 0.80 |
| CLIP [41] | ViT-B/16 | VLM | WIT-400M | 9.02 | 14.19 | 8.44 | 0.50 |
| MiDas [42] | ViT-L/16 | Depth | MIX-6 | 10.65 | 17.42 | 9.87 | 0.69 |
| DINOv2 [38] | ViT-B/14 | SSL | LVD-142M | 16.05 | 24.27 | 14.97 | 2.40 |
@@ 5.3 Analysis
Fig. 6 reports our model’s AP score as a function of the training data size. This underscores the critical importance of dataset size in open vocabulary monocular 3D object detection tasks and suggests that our model may achieve even better performance with more extensive training data.
@@
| Synthetic+Real | 175k | 24.77 | 16.05 |
5.4 Zero-Shot Generalization Performance
Fig. 7 presents OVMono3D-LIFT prediction on in-the-wild images from COCO [31]. We visualize with intrinsics of , , , where is the input image resolution. The results show 2D projections of our predictions are well-aligned with the objects, and their top-down views closely match the visual scene layout. Even on completely out-of-distribution categories such as elephant and apple, our method produces promising results. This suggests that our model demonstrates zero-shot generalization ability on real-world images.
6 Discussion
We investigate OVMono3D, an under-explored task of recognizing and localizing objects from any categories in 3D using a single image. We identify unique challenges, notably data scarcity and limitations of standard evaluation metrics. We propose simple yet effective approaches, including geometry-based OVMono3D-GEO and learning-based OVMono3D-LIFT — both decouple 2D detection from 3D bounding box prediction. Such design enables the full utilization of off-the-shelf open-vocabulary 2D detectors pre-trained on large-scale datasets. Our analysis pinpoints the key components in the framework, including pre-trained image encoder, base 2D detectors, and the impact of dataset scaling. We further demonstrate the zero-shot generalizability of our approach on in-the-wild images. Our findings suggest that dataset scale and accurate depth perception remain the major bottlenecks in this task. One promising direction may be to develop unsupervised learning that harness the abundance of unlabeled images. We hope this work inspires future research toward advancing this task.
Acknowledgement.
The authors acknowledge Research Computing at The University of Virginia for providing computational resources and technical support that have contributed to the results reported within this publication.
Supplementary Material
Sec. A presents per-category performance on novel classes in 2D and 3D for OVMono3D-GEO and OVMono3D-LIFT. Sec. B provides additional qualitative visualizations on various datasets and compares predictions between OVMono3D-GEO and OVMono3D-LIFT. Sec. C discusses failure cases, highlighting challenges such as occlusions, out-of-distribution objects, and small or distant instances.
Appendix A Per-category Performance on novel classes
We show per-category performance on Average Precision (AP) and Average Recall (AR) on both 2D and 3D for OVMono3D-GEO and OVMono3D-LIFT in Tabs. 8 and 9, respectively.
| Category | AP | AP | AR | AR |
| Board | 29.41 | 2.06 | 34.05 | 7.78 |
| Printer | 31.05 | 8.96 | 41.77 | 22.99 |
| Painting | 33.39 | 2.09 | 44.80 | 9.04 |
| Microwave | 31.60 | 14.06 | 39.45 | 28.27 |
| Tray | 15.65 | 3.40 | 22.30 | 13.11 |
| Podium | 46.66 | 36.71 | 53.85 | 46.92 |
| Cart | 31.49 | 32.52 | 38.25 | 41.75 |
| Tram | 4.72 | 0.61 | 6.91 | 1.58 |
| Easy Categories | 28.00 | 12.55 | 35.17 | 21.43 |
| Monitor | 18.84 | 6.55 | 30.02 | 19.36 |
| Bag | 13.98 | 8.20 | 23.46 | 21.45 |
| Dresser | 27.43 | 14.87 | 38.22 | 30.86 |
| Keyboard | 12.06 | 7.23 | 25.41 | 18.56 |
| Drawers | 19.87 | 9.29 | 32.67 | 21.42 |
| Computer | 15.73 | 5.73 | 25.00 | 19.70 |
| Kitchen Pan | 5.56 | 7.36 | 14.22 | 16.67 |
| Potted Plant | 15.05 | 10.65 | 28.06 | 27.10 |
| Tissues | 5.88 | 3.04 | 11.55 | 9.86 |
| Rack | 9.19 | 3.54 | 12.50 | 8.17 |
| Toys | 10.35 | 8.84 | 21.94 | 19.17 |
| Phone | 3.30 | 5.00 | 9.52 | 12.62 |
| Soundsystem | 13.09 | 7.43 | 20.47 | 19.53 |
| Fireplace | 17.48 | 0.42 | 39.51 | 3.86 |
| Hard Categories | 13.42 | 7.01 | 23.75 | 17.74 |
| All Categories | 18.72 | 9.03 | 27.91 | 19.08 |
| Category | AP | AP | AR | AR |
| Board | 38.90 | 4.83 | 54.51 | 20.52 |
| Printer | 31.02 | 16.23 | 44.08 | 43.95 |
| Painting | 34.99 | 2.80 | 50.23 | 19.77 |
| Microwave | 32.76 | 30.31 | 42.91 | 56.27 |
| Tray | 17.75 | 10.11 | 32.79 | 30.16 |
| Podium | 45.63 | 48.37 | 57.18 | 73.08 |
| Cart | 37.54 | 47.31 | 50.75 | 71.50 |
| Tram | 22.86 | 4.71 | 34.00 | 9.15 |
| Easy Categories | 32.68 | 20.58 | 45.81 | 40.55 |
| Monitor | 21.46 | 9.44 | 38.64 | 35.44 |
| Bag | 15.90 | 15.61 | 30.22 | 36.42 |
| Dresser | 30.54 | 29.08 | 48.49 | 60.26 |
| Keyboard | 14.60 | 9.13 | 36.71 | 25.07 |
| Drawers | 21.48 | 43.04 | 43.00 | 68.33 |
| Computer | 16.80 | 7.44 | 35.30 | 28.56 |
| Kitchen Pan | 6.81 | 9.98 | 21.57 | 35.10 |
| Potted Plant | 12.94 | 6.66 | 33.23 | 24.62 |
| Tissues | 8.81 | 12.45 | 21.55 | 32.68 |
| Rack | 13.02 | 10.21 | 27.33 | 35.33 |
| Toys | 10.91 | 5.24 | 29.31 | 22.08 |
| Phone | 6.23 | 3.89 | 22.14 | 16.43 |
| Soundsystem | 14.52 | 13.22 | 31.86 | 31.86 |
| Fireplace | 19.50 | 13.16 | 47.62 | 34.06 |
| Hard Categories | 15.25 | 13.47 | 33.36 | 34.73 |
| All Categories | 21.59 | 16.05 | 37.88 | 36.85 |
Appendix B More Qualitative Results
Additional qualitative visualizations of OVMono3D-LIFT are provided for Omni3D [5] outdoor, indoor subsets, and COCO [31] in-the-wild images in Figs. 8, 9 and 10, respectively.
Appendix C Failure Cases
Fig. 12 shows failure cases of OVMono3D-LIFT on COCO [31] images. In Fig. 12a, the relative position from a top-down view is incorrect, indicating that our model sometimes predicts the wrong object depth. In Fig. 12b, our model fails to predict the correct size and pose for the bear, suggesting that it may struggle with totally out-of-distribution objects. In Fig. 12c, our model fails to detect the person and bus in the distance, indicating that it may not perform well on small and distant objects. In Fig. 12d, our method fails to identify the mirror and incorrectly detects the object in the mirror. These failure cases suggest that our model still has room for improvement. Future research could explore better model architectures and weakly supervised learning techniques to address these shortcomings.

### v2

@@ 5.1 Experimental Setup
The experiments are conducted using the Omni3D [brazil2023omni3d] dataset, the largest dataset for monocular 3D object detection across both indoor and outdoor scenes. Omni3D is repurposed and combined from six established sources including SUN RGB-D [song2015sun], ARKitScenes [baruch2021arkitscenes], Objectron [ahmadyan2021objectron], Hypersim [roberts2021hypersim], nuScenes [caesar2020nuscenes], and KITTI [geiger2012we]. It comprises a substantial 234k images, 3 million labeled 3D bounding boxes, and covers 98 distinct object categories.
During training, we use the 50 categories Cube R-CNN [brazil2023omni3d] trained on as the base categories. For evaluation, we select 22 categories from the remaining classes as novel categories. These categories are chosen based on two criteria: the number of test instances and the precision of category naming. To facilitate a detailed assessment of zero-shot generalization capabilities, we further divide these categories into easy and hard subsets according to object visibility. See Appendix A for detailed splits.
Besides evaluating novel categories in Omni3D [brazil2023omni3d], we perform zero-shot evaluation on a novel dataset with novel camera models. CityScapes3D [gahlert2020cityscapes] includes base categories from the self-driving domain. We use the same evaluation split as DetAny3D [zhang2025detect].
Baselines.
For novel categories, we compare against OVM3D-Det [huang2024training], which generates pseudo 3D labels by combining Grounding DINO [liu2023grounding], SAM [kirillov2023segment], UniDepth [piccinelli2024unidepth], and LLM-generated size priors. We evaluate their labeling pipeline as a baseline detector against ground truth annotations. We also include comparisons with DetAny3D [zhang2025detect]. For base categories, we compare against closed-vocabulary methods including SMOKE [liu2020smoke], FCOS3D [wang2021fcos3dfullyconvolutionalonestage], PGD [wang2022probabilistic], ImVoxelNet [rukhovich2022imvoxelnet], Cube R-CNN [brazil2023omni3d], and UniMODE [li2024unified3dobjectdetection], as well as open-vocabulary methods including OVM3D-Det [huang2024training] and DetAny3D [zhang2025detect].
Evaluation Metrics.
We report the mean Average Precision in 3D (AP) computed across all evaluated categories. Following the Omni3D [brazil2023omni3d] evaluation protocol, predictions are matched to ground truth using 3D Intersection over Union (IoU), which measures the volumetric overlap between predicted and ground truth 3D bounding boxes. The evaluation is conducted across IoU thresholds , with the final AP representing the mean average precision across all thresholds and categories. For evaluation on novel categories and novel datasets, we additionally report target-aware metrics (see Sec. 4).
@@
Our implementation is based on PyTorch3D [ravi2020accelerating] and Detectron2 [wu2019detectron2]. We initialize the image feature encoder with pre-trained DINOv2 base weights [oquab2023dinov2]. The model is trained on eight NVIDIA A100 GPUs for five days. For reference, DetAny3D [zhang2025detect] is trained on 64 A100 GPUs for two weeks. The models are trained for 116k steps with a batch size of 64 using SGD optimizer with an initial learning rate of 0.0012, which decays by a factor of 10 at 60% and 80% of training. We apply standard image augmentations including random horizontal flipping and resizing during training.
Due to computational constraints, models trained in the analysis section ( Sec. 5.3) employ a resource-efficient setting where the pre-trained image encoder is kept frozen during training and the depth estimator is excluded. Unless explicitly stated otherwise, all methods use Grounding DINO [liu2023grounding] for 2D detection on novel categories, and all methods requiring pseudo depth input use UniDepthv2 [piccinelli2025unidepthv2].
@@ 5.2 Model Performance
Fig. 5 shows qualitative results on the Omni3D test set. In comparison with Cube R-CNN [brazil2023omni3d], OVMono3D-LIFT detects not only objects of base categories, but also novel categories that are unseen at the training time.
Novel Category and Dataset Performance.
Tab. 1 shows that OVMono3D-LIFT achieves state-of-the-art performance on novel categories and novel datasets.
Geometry-based methods OVM3D-Det [huang2024training] and OVMono3D-GEO show lower performance, particularly on Cityscapes3D [gahlert2020cityscapes], due to their sensitivity to occlusions and noisy depth estimation. The performance improvement from LIFT* to LIFT underscores the importance of incorporating additional geometric information from foundation depth estimators into the model. This enhanced performance demonstrates that our geometry-informed 2D lifting design effectively exploits 2D data-driven priors, including OV 2D detectors, self-supervised features, and metric depth estimators. By decomposing the task into OV 2D detection and 2D-to-3D lifting, our approach addresses challenges associated with low-quality and limited 3D annotations.
@@
Tab. 2 compares OVMono3D-LIFT with baselines on base categories. OVMono3D-LIFT achieves the best performance among both closed-vocabulary and open-vocabulary methods, with an overall AP of 29.63. It surpasses Cube R-CNN [brazil2023omni3d] by 5.95 points and UniMODE [li2024unified3dobjectdetection] by 1.43 points, while additionally providing generalization to novel categories. Compared to the open-vocabulary baseline DetAny3D [zhang2025detect], it achieves a 4.71-point improvement, demonstrating strong performance on base categories. Notably, OVMono3D-LIFT* also achieves competitive results with 27.09 overall AP, ranking second among open-vocabulary methods. These results validate the effectiveness of our approach.
5.3 Analysis
| Methods | IoU↑ | XY↑ | Depth↑ | Size↑ | Pose↑ |
| OVMono3D-GEO | 18.69 | 47.96 | 30.97 | 18.68 | 37.84 |
| OVMono3D-LIFT* | 18.11 | 49.39 | 27.23 | 48.10 | 72.44 |
| OVMono3D-LIFT | 22.03 | 52.91 | 32.59 | 49.64 | 73.57 |
Disentangled Metrics.
To analyze the contributions of different predicted attributes to 3D bounding box errors, we report disentangled IoU for position (XY), depth, dimensions, and pose. For each attribute, we compute the IoU of a modified 3D bounding box where only that attribute is predicted, while all others are set to their ground-truth values. The IoU is then computed against the ground-truth cube. To facilitate the comparison, for all methods we use ground truth 2D box as input.
Tab. 3 exhibits the overall and disentangled IoU for different attribute predictions. For LIFT* and LIFT methods, object depth prediction consistently contributes the most error to the overall prediction. This indicates that object depth estimation is the primary bottleneck of our task. Furthermore, the size and pose predictions in GEO method exhibit larger errors, indicating that geometry-based methods are not effective for these predictions. This underscores the necessity of developing learning-based methods.
2D Bounding Box Input.
Tab. 4 evaluates the impact of 2D detectors on both 2D and 3D detection performance for novel categories. We test two state-of-the-art 2D detectors: YOLO-World [cheng2024yolo] and Grounding DINO [liu2023grounding]. Grounding DINO consistently outperforms YOLO-World across all evaluation metrics, demonstrating that Grounding DINO is a preferable off-the-shelf 2D detector for OVMono3D.
@@ 5.3 Analysis
| YOLO-World [cheng2024yolo] | 19.99 | 20.31 | 31.50 | 19.42 | 2.85 |
| Grounding DINO [liu2023grounding] | 21.44 | 24.08 | 37.45 | 23.24 | 4.00 |
@@
Fig. 6 reports our model’s AP score as a function of the training data size. This underscores the critical importance of dataset size in OVMono3D tasks and suggests that our model may achieve even better performance with more extensive training data.
Pre-trained feature extractor Selection.
Tab. 5 shows the impact of pretrained feature extractors on 3D detection performance for novel categories. DINOv2 achieves the best performance across all evaluation metrics. This result underscores the effectiveness of DINOv2’s representations for the 3D detection task. Our findings align with recent studies that highlight DINOv2’s strong zero-shot capabilities in understanding depth, multi-view correspondences, and relative pose [krishnan2025omninocs, goodwin2022zero, zhang2024tale], indicating that DINOv2’s 3D-aware image features are highly suitable for this task.
For a more detailed analysis on the role of synthetic data and quantitative evaluation of naming ambiguity issues in current benchmarks, refer to Appendix D.
| Feat. Extractor | Supervision | AP↑ | AP↑ | AP↑ | AP↑ |
| MAE [he2022masked] | SSL | 7.72 | 11.95 | 6.43 | 0.80 |
| CLIP [radford2021learning] | VLM | 9.02 | 14.19 | 8.44 | 0.50 |
| MiDas [ranftl2020towards] | Depth | 10.65 | 17.42 | 9.87 | 0.69 |
| DINOv2 [oquab2023dinov2] | SSL | 16.04 | 24.57 | 14.97 | 2.21 |
5.4 Zero-Shot Generalization Performance
LABEL:fig:coco presents OVMono3D-LIFT prediction on in-the-wild images from COCO [lin2014microsoft]. The results show 2D projections of our predictions are well-aligned with the objects, and their top-down views closely match the visual scene layout. Even on completely out-of-distribution categories such as elephant and apple, our method produces promising results. This suggests that our model demonstrates zero-shot generalization ability on real-world images. See Fig. 9 in Supplementary Material for more qualitative results on COCO.
6 Discussion
We investigate OVMono3D, an under-explored task of recognizing and localizing objects from any categories in 3D using a single image. We identify unique challenges, notably data scarcity and limitations of standard evaluation metrics. We propose simple yet effective approaches, including geometry-based OVMono3D-GEO and learning-based OVMono3D-LIFT — both decouple 2D detection from 3D bounding box prediction. Such design enables the full utilization of off-the-shelf open-vocabulary 2D detectors pre-trained on large-scale datasets. Our analysis pinpoints the key components in the framework, including 3D-aware image features, base 2D detectors, and the impact of dataset scaling. We further demonstrate the zero-shot generalizability of our approach on in-the-wild images. Our findings suggest that dataset scale and accurate depth perception remain the major bottlenecks in this task. One promising direction may be to develop unsupervised learning that harness the abundance of unlabeled images. We hope this work inspires future research toward advancing this task.
7 Limitations
Due to the lack of 3D detection ground truth labels in COCO [lin2014microsoft], the qualitative zero-shot evaluation is not feasible to perform. Additionally, our method requires accurate camera intrinsics as input; however, for in-the-wild images, the estimated intrinsics can be inaccurate, leading to errors in prediction. Furthermore, the use of computationally heavy components, such as Grounding DINO [liu2023grounding] and DINOv2 [oquab2023dinov2], results in slower inference speed compared to Cube R-CNN [brazil2023omni3d], which should be addressed in future work. See Appendix C for visualizations of failure cases.
8 Acknowledgement
The authors acknowledge the University of Virginia Research Computing and Data Analytics Center, Advanced Micro Devices AI and HPC Cluster Program, Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, and National Artificial Intelligence Research Resource (NAIRR) Pilot for computational resources, including the Anvil supercomputer (National Science Foundation award OAC 2005632) at Purdue University and the Delta and DeltaAI advanced computing resources (National Science Foundation award OAC 2005572). This work was partly supported by an Adobe Research Gift.
Supplementary Material
Sec. A presents per-category performance on novel classes for OVMono3D-GEO and OVMono3D-LIFT. Sec. B provides additional qualitative visualizations on various datasets and compares predictions between OVMono3D-GEO and OVMono3D-LIFT. Sec. C discusses failure cases, highlighting challenges such as occlusions, out-of-distribution objects, and small or distant instances. Sec. D provides more analysis on synthetic data and the naming ambiguity issue in current benchmarks.
Appendix A Per-category Performance on novel classes
We show per-category performance on 3D Average Precision (AP) for OVMono3D-GEO and OVMono3D-LIFT in Tab. 6.
| Category | GEO | LIFT |
| Board | 12.42 | 9.92 |
| Printer | 34.03 | 35.90 |
| Painting | 5.16 | 6.31 |
| Microwave | 33.47 | 48.54 |
| Tray | 10.35 | 15.56 |
| Podium | 41.74 | 62.18 |
| Cart | 32.22 | 54.00 |
| Tram | 0.25 | 8.65 |
| Easy Categories | 21.20 | 30.13 |
| Monitor | 18.25 | 13.92 |
| Bag | 25.48 | 23.96 |
| Dresser | 29.78 | 36.71 |
| Keyboard | 15.44 | 12.58 |
| Drawers | 30.01 | 57.21 |
| Computer | 13.14 | 13.77 |
| Kitchen Pan | 15.44 | 19.90 |
| Potted Plant | 20.13 | 6.07 |
| Tissues | 13.49 | 18.28 |
| Rack | 14.60 | 15.74 |
| Toys | 24.07 | 21.70 |
| Phone | 22.21 | 11.37 |
| Soundsystem | 17.73 | 17.69 |
| Fireplace | 24.44 | 19.76 |
| Hard Categories | 20.30 | 20.62 |
| All Categories | 20.63 | 24.08 |
Appendix B More Qualitative Results
Additional qualitative visualizations of OVMono3D-LIFT are provided for Omni3D [brazil2023omni3d] outdoor, indoor subsets, and COCO [lin2014microsoft] in-the-wild images in Figs. 7, 8 and 9, respectively. For COCO images, we visualize with intrinsics of , , , where is the input image resolution.
Appendix C Failure Cases
Fig. 11 shows failure cases of OVMono3D-LIFT on COCO [lin2014microsoft] images. In Fig. 11a, the relative position from a top-down view is incorrect, indicating that our model sometimes predicts the wrong object depth. In Fig. 11b, our model fails to predict the correct size and pose for the bear, suggesting that it may struggle with totally out-of-distribution objects. In Fig. 11c, our model fails to detect the person and bus in the distance, indicating that it may not perform well on small and distant objects. In Fig. 11d, our method fails to identify the mirror and incorrectly detects the object in the mirror. These failure cases suggest that our model still has room for improvement. Future research could explore better model architectures and weakly supervised learning techniques to address these shortcomings.
Appendix D More Analysis
@@
We conducted an ablation study using synthetic data for OVMono3D-LIFT under resource-constrained conditions, with a frozen image encoder and excluded depth estimator. The synthetic data comes from Hypersim [roberts2021hypersim], which provides indoor images rendered from artist-created meshes and serves as the sole synthetic data source in Omni3D [brazil2023omni3d].
@@
| Synthetic+Real | 175k | 24.77 | 16.04 |
Naming Ambiguity issue.
We quantitatively evaluate naming ambiguity in current 3D benchmarks using SUN RGB-D [song2015sun] as an example. For each object instance, we cropped its 2D bounding box and computed CLIP similarity scores between the visual features and all category text embeddings. We then aggregated these similarity vectors by ground-truth category and computed average similarities to form a confusion matrix, applying softmax normalization.
As shown in Fig. 12, SUN RGB-D annotations exhibit weaker self-correlation than COCO [lin2014microsoft], indicating less distinct category boundaries. This reflects SUN RGB-D’s highly similar category names (e.g., ”table” vs. ”desk”). In open-vocabulary settings, such similarity creates false negatives when models correctly identify a table as a desk—a distinction often acceptable in real-world applications. Therefore, our proposed target-aware evaluation is essential for datasets with ambiguous category definitions, unlike the well-differentiated categories in COCO.